40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Mar 18 12:22:04 2011, kiwamu, Update, Green Locking, Y arm plan for today 
    Reply  Fri Mar 18 16:06:30 2011, kiwamu, Update, Green Locking, Re: Y arm plan for today 
       Reply  Fri Mar 18 16:31:11 2011, Suresh , Update, Green Locking, Re: Y arm plan for today 
Message ID: 4411     Entry time: Fri Mar 18 12:22:04 2011     Reply to this: 4413
Author: kiwamu 
Type: Update 
Category: Green Locking 
Subject: Y arm plan for today 

 Prior to the works on the Y end setup I propose to perform the temperature scan business like Koji and Suresh did before (see this entry).

This business will allow us to easily find a beatnote at 532nm after the installation on the Y end.

 I guess the right persons for this work are Bryan and Suresh.

Bryan will have a safety guidance from Steve in this after noon. So after that they can start working on it.

 

/* - - - coarse plan - - - */

* remove Alberto's laser from the AS table

* setup Alberto's laser on the PSL table

* put some stuff such as lenses, mirrors and etc. (Use the IR beam picked off after the doubling crystal for the main laser source)

* mode matching

* measurement

 

Which laser are we going to use,  Alberto's laser or MOPA laser ?

ELOG V3.1.3-